# Transforming the Bootstrap: Using Transformers to Compute Scattering Amplitudes in Planar $\mathcal{N}=4$ Super Yang-Mills Theory

### v1

@@ Abstract
**footnotetext: Denotes equal contribution
  • May 2024
@@ 1 Introduction
Figure 1: Sample Feynman diagrams for the process $gg\rightarrow Hg$ at two loops (left) and eight loops (right).
A recently-developed alternative technique, known as the amplitude bootstrap [11, 12, 13], attempts to directly construct candidate solutions for multi-loop amplitudes. It has mainly been applied so far to a simpler relative of QCD, called planar ${\cal N}=4$ super-Yang-Mills theory (SYM). The amplitude bootstrap leverages the rich, yet highly constrained, analytical structure of amplitudes that arises from the particular recurrent features of the integrals involved. Using this technique, the form of the amplitude can be determined a priori, and the finite-dimensional solution space at a given loop order can be strongly constrained through a large system of linear relations with integer coefficients. Many of the linear relations are found by analyzing the lower-loop results. A small number of additional constraints, derived from behavior in physical limits, can then be applied in order to obtain a unique solution.
The amplitude bootstrap allows for the computation of amplitudes up to eight loops in SYM [14], vs. two loops using traditional Feynman diagram methods for the same quantity in QCD [15], as depicted in Figure 1. However, the number of linear equations to be solved when using the bootstrap technique scales exponentially with loop order, making it intractable beyond eight loops. This situation necessitates the development of new methods that may exploit hitherto unobserved patterns in the data in order to simplify the computation of scattering amplitudes at higher loops.
Notably, amplitudes in SYM can be expressed as sets of tokenizable “words” with integer coefficients; see Section 2. We refer to these words as keys which index into the coefficients, and to a key-coefficient pair as an element. Exploiting the many linear relations between these elements amounts to solving a large system of linear equations, where the integer solutions are hard to discover, but easy to verify. Since both the keys and the integers can be represented as sequences of tokens, we can train deep learning models such as Transformers to predict the coefficient associated with each key. In principle, one can easily verify any model predictions using the known linear relations.
@@ 1.1 Related Work
The amplitude bootstrap program has a long tradition, dating back to [11]. A recent review can be found in [22], and the specific data we use is from [14]. Our work supplements the traditional approach by offering a novel problem-solving framework, where human intelligence is augmented by artificial intelligence to further push the state-of-the-art for amplitude calculations.
@@ 2.1 Symbols: A Simple Language for Amplitudes
Here $\mathcal{L}_{m}=\{l_{1},\dots,l_{m}\}$ is the symbol alphabet containing $m$ letters $l_{i}$, which are in turn functions of the particles’ four-momenta, and $C^{l_{i_{1}},\ldots,l_{i_{2L}}}$ is a ${2L}$-fold tensor of integer coefficients, most of which are zero. In other words, a solution for the symbol at loop $L$ can be represented by $m^{2L}$ integers, with each sequence of $2L$ letters (i.e., $l_{i_{1}}\otimes\cdots\otimes l_{i_{2L}}$) serving as a key indexing into the integer-valued tensor $C^{l_{i_{1}},\ldots,l_{i_{2L}}}$. More details about the map from generalized polylogarithms to symbols are given in A.
@@ 2.2 Linear Relationships among Symbol Elements
A concrete instance of the relation (6) at five loops is
@@ 3 Implementation Details
In most experiments, we use encoder-decoder Transformers, which contain a bidirectional Transformer encoder and an autoregressive Transformer decoder linked by a cross-attention mechanism [16]. Both encoder and decoder have the same number of layers (up to $8$), the same number of attention heads ($8$ or $16$), and the same dimension ($d=256$, $d=512$ or $d=1024$). Henceforth, we describe a model with $N$ layers in the encoder and $N$ layers in the decoder as an $N$-layer Transformer.
@@
In the default model, we use a learnable positional encoding in both the encoder and decoder [16]; alternative schemes are explored in C. The optimizer is Adam [41], with a learning rate of $10^{-4}$ and a flat learning rate schedule. All models are implemented in PyTorch [42] and trained on a single NVIDIA V100 GPU with $32$ GB of memory, or on larger architectures (A100).
Throughout the paper, we define an epoch as a pass over 300,000 key-coefficient pairs, as opposed to the more common definition of one full pass over all training data. This makes the notion of epoch size more comparable between different experiments that employ different amounts of data at different loops. At the end of each epoch, the model is evaluated on a held-out test set.
@@ 5 Quad Representation of Symbols
At higher loops, a more compact representation of the form factor symbol is necessary for efficient training. For example, the loop $L=7$ symbol has $93$ million nonzero-coefficient elements, almost 19 times as many as the $L=6$ symbol. Compressing the data using the quad representation can significantly improve the model training speed by reducing the number of nonzero-coefficient elements to a more manageable $7.3$ million.
@@ 9 Conclusions and Future Work
We thank Yang-Hui He, Romuald Janik, Matt Schwartz, Jesse Thaler, Dmitris Papailiopoulos, Jordan Ellenberg, Gary Shiu, and Yiqiao Zhong for fruitful discussions. We would like to thank Phil Wang (“lucidrains”) for use of his implementation of the RoPE and xPos relative position encodings. The work was supported in part by the U.S. Department of Energy (DOE) under Award No. DE-FOA-0002705, KA/OR55/22 (AIHEP). LD and TC are additionally supported by the U.S. Department of Energy Award No. DE-AC02-76SF00515. MW was supported by the research grant 00025445 from Villum Fonden.
@@ Appendix C Coefficient-from-Key: Model Ablations
We perform a number of ablations in order to better understand the model performance for the coefficient-from-key experiment in Section 4, with an eye towards building a future model that can generalize across loops. Here we consider several different positional encoding schemes, explore encoder-only and decoder-only architectures, and vary both the number of layers in the model and the dimension $d$ of the encoded tokens. See Tables 5 and 6 for an overview of the results.
@@
C.1 Architecture Ablations: Relative Position Encoding, Encoder-only, Decoder-only

### v2

@@ Abstract
SLAC-PUB-17774
**footnotetext: Denotes equal contribution**footnotetext: Authors to whom correspondence should be addressed
  • tianji@slac.stanford.edu, garrett.merz@wisc.edu, fcharton@meta.com, nolte@meta.com, matthias.wilhelm@nbi.ku.dk, kyle.cranmer@wisc.edu, lance@slac.stanford.edu
  • Sept 2024
@@ 1 Introduction
Figure 1: Sample Feynman diagrams for the process $gg\rightarrow Hg$ at two loops (left) and eight loops (right) in QCD. The same diagrams contribute in SYM, where the Higgs boson $H$ and top quark ($t$) triangle is replaced by a particular local operator in the theory, and the process is referred to as a form factor.
A recently-developed alternative technique, known as the amplitude bootstrap [11, 12, 13], attempts to directly construct candidate solutions for multi-loop amplitudes. It has mainly been applied so far to a simpler relative of QCD, called planar ${\cal N}=4$ super-Yang-Mills theory (SYM). The amplitude bootstrap circumvents many of the computational and numerical challenges that arise from the Feynman diagram approach. It leverages the rich, yet highly constrained, analytical structure of amplitudes that arises from the particular recurrent features of the integrals involved. Using this technique, the form of the amplitude can be determined a priori, and the finite-dimensional solution space at a given loop order can be strongly constrained through a large system of linear relations with integer coefficients. Many of the linear relations are found by analyzing the lower-loop results. A small number of additional constraints, derived from behavior in physical limits, can then be applied in order to obtain a unique solution.
The amplitude bootstrap allows for the computation of amplitudes up to eight loops in SYM [14], vs. two loops using traditional Feynman diagram methods for the same quantity in QCD [15], as depicted in Figure 1. However, when using the bootstrap technique, the number of linear relations to be solved and the number of unknown coefficients both increase by a factor of about 4 at each subsequent loop order [14]. Since it is not possible to determine the minimal set of independent equations in advance, the number of relations can be several times larger than the number of unknown coefficients. The computational cost of generating and solving the equations thus increases by at least a factor of $164^{2}=$ 16 at each loop order. Performing computations by this method becomes infeasible beyond about loop $L=9$. This situation necessitates the development of new methods that may exploit hitherto unobserved patterns in the data in order to simplify the computation of scattering amplitudes at higher loops.
Notably, amplitudes in SYM can be expressed as sets of tokenizable “words” with integer coefficients; see Section 2. We refer to these words as keys which index into the coefficients, and to a key-coefficient pair as an element. Exploiting the many linear relations between these elements amounts to solving a large system of linear equations (as explained above), where the integer solutions are hard to discover, but easy to verify.
In our case, verification includes first checking that the proposed expression is a legitimate function in the appropriate space (see A), and then comparing its behavior in a physical limit to the predictions of the form-factor operator product expansion (FFOPE) [16, 17]. The latter constraint is stringent enough to fix the answer uniquely [14].
Since both the keys and the integers can be represented as sequences of tokens, we can train deep learning models such as Transformers to predict the coefficient associated with each key.
@@
We emphasize that the goal of these experiments is not to optimize performance. Rather it is to show that properties of amplitudes can be learned by AI models, and to probe the learning dynamics.
@@ 1.1 Related Work
The amplitude bootstrap program has a long tradition, dating back to ref.  [11]. A recent review can be found in ref.  [24], and the specific data we use is from ref.  [14]. Our work supplements the traditional approach by offering a novel problem-solving framework, where human intelligence is augmented by artificial intelligence to further push the state-of-the-art for amplitude calculations.
@@ 2.1 Symbols: A Simple Language for Amplitudes
Here $\mathcal{L}_{m}=\{l_{1},\dots,l_{m}\}$ is the symbol alphabet containing $m$ letters $l_{i}$, which are in turn functions of the particles’ four-momenta, and $C^{l_{i_{1}},\ldots,l_{i_{2L}}}$ is a ${2L}$-fold tensor of integer coefficients, most of which are zero. In other words, a solution for the symbol at loop $L$ can be represented by $m^{2L}$ integers, with each sequence of $2L$ letters (i.e., $l_{i_{1}}\otimes\cdots\otimes l_{i_{2L}}$) serving as a key indexing into the integer-valued tensor $C^{l_{i_{1}},\ldots,l_{i_{2L}}}$. Symbol terms encode information about the derivatives and discontinuities of the polylogarithms, and do not correspond directly to individual Feynman diagrams. More details about the map from generalized polylogarithms to symbols are given in A.
@@
One can recover the iterated integral representation from the symbol, up to constants that can be recovered in a similar fashion. For the one-loop form factor given in eq. (4), its iterated integral representation is
$\mathcal{F}_{\tt 3gFF}^{(1)}=2\,\big{[}{\rm Li}_{2}(1-bc)+{\rm Li}_{2}(1-ca)+{\rm Li}_{2}(1-ab)\big{]}\,,$ (5)
[…]
@@ 2.2 Linear Relationships among Symbol Elements
A concrete instance of relation (7) at five loops is
@@ 3 Implementation Details
In most experiments, we use encoder-decoder Transformers, which contain a bidirectional Transformer encoder and an autoregressive Transformer decoder linked by a cross-attention mechanism [18]. Both encoder and decoder have the same number of layers (up to $8$), the same number of attention heads ($8$ or $16$), and the same dimension ($d=256$, $d=512$ or $d=1024$). For all models, the tokenizer dimension and Transformer dimension $d$ are the same (i.e., $256$, $512$, or $1024$). Henceforth, we describe a model with $N$ layers in the encoder and $N$ layers in the decoder as an $N$-layer Transformer.
@@
In the default model, we use a learnable positional encoding in both the encoder and decoder [18]; alternative schemes are explored in C. The optimizer is Adam [43], with a learning rate of $10^{-4}$ and a flat learning rate schedule. We do not observe significant performance improvements from changes to the learning rate schedule or addition of learning-rate warmup. We believe this is due to the small size of the Transformers we use: in our experience, warmup becomes more useful for larger models.
All models are implemented in PyTorch [44] and trained on a single NVIDIA V100 GPU with $32$ GB of memory, or on larger architectures (A100).
Throughout the paper, we define an epoch as a pass over 300,000 key-coefficient pairs, as opposed to the more common definition of one full pass over all training data. This makes the notion of epoch size more comparable between different experiments that employ different amounts of data at different loops. At the end of each epoch, the model is evaluated on a held-out test set. The train-test split is performed randomly. In all experiments, metrics such as accuracy are quoted on the held-out test set, not the full symbol. For example, for a training set of 100,000 elements and a test set of 1,000 elements, an accuracy of 90% refers to predicting 900 of those 1,000 elements correctly.
Accuracies are measured on a test set of $N_{test}$ randomly sampled elements. The accuracy for each element is a binary variable (correct $=1$, wrong $=0$), so we may assume that the errors are binomial. Applying the central limit theorem to the average, we have a 95% confidence interval of $1.96\times\sqrt{a(1-a)/N_{test}}\approx 2\sqrt{(1-a)/N_{test}}$ due to statistical error, for an accuracy $a\approx 1$. For $a=99\%$ and a test set size of $N_{test}=10,000$, this interval is $0.2\%$. There is also uncertainty in the accuracy due to the model’s initialization.
Because the goal of these experiments is not to optimize final accuracy, we do not perform any fine-grained hyperparameter scans. For this reason, we do not typically have a separate validation set in addition to the test set. However, in order to determine whether the specific choice of training and test set might play a role, in C.2 we repeat one of our experiments for a few different train/test splits using the same set of hyperparameters chosen for the main work. We find qualitatively similar performance.
@@ 5 Quad Representation of Symbols
At higher loops, a more compact representation of the form factor symbol is necessary for efficient training. For example, the loop $L=7$ symbol has $93$ million nonzero-coefficient elements, almost 19 times as many as the $L=6$ symbol. How long might it take to train on such a large symbol? At loop $L=5$, performing the nonzero coefficient-from-key prediction task to $>99\%$ accuracy takes about 22 passes through the training set and 0.7 hours. At loop $L=6$, the same task takes 11 passes through the training set and over 54 hours. The time to reach $>99\%$ accuracy scales at least linearly with the number of elements. If we assume this scaling continues to $L=7$, it will take at least $54\times 19\approx 1000$ hours, or 43 days.
Compressing the data using the quad representation can significantly improve the model training speed by reducing the number of nonzero-coefficient elements to a more manageable $7.3$ million. Under the linear scaling paradigm, we would predict a more manageable 80 hours, or under 4 days.
@@
The Transformer’s ability to learn the sign appears to be governed largely by model capacity and training dataset size. Thus, in regions where the model sees enough data and is large enough to consistently learn the sign, accuracy is consistently $>90\%$, while in regions where the model is too small, the sign is not learned, and accuracy stays at or below $50\%$. When we are near a model capacity threshold (for this task, $4$ layers and $d=512$) the model sometimes learns the sign and sometimes does not.
@@ 6 Model Characterization via Relationship Accuracy
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ 9 Conclusions and Future Work
Transformers are well-studied, but they are not the only architecture with the potential to learn the structure of scattering amplitudes. In C.1, we train a Long-Short-Term Memory recurrent neural network (LSTM) [45] and a Gated Recurrent Unit (GRU) [46] on the $L=5$ symbol. In both cases, we are able to attain comparable performance to a Transformer. It may be interesting to explore other types of architectures as well.
@@
The methods in this paper can be applied straightforwardly to other SYM problems where there is multi-loop data encoded by symbols. Additional challenges arise when considering a similar approach to amplitudes in QCD. While many of the generalized polylogarithms appearing in SYM also appear in QCD, the QCD result does not have uniform weight $2L$, but generally has all weights from $2L$ down to zero (rational functions). The lower-weight terms typically have complicated prefactors that are rational functions of the kinematic variables. Also, the behavior in physical limits is not understood as well in QCD as in SYM. A better understanding of how to organize and bootstrap QCD amplitudes would clearly benefit any machine-learning approaches to determining them.
@@
The data that support the findings of this study are openly available at the following URL/DOI: https://10.5281/zenodo.11218272 [47]. A prior version of this work was presented in a non-archival form at the 2023 NeurIPS Machine Learning for the Physical Sciences workshop [48]. We thank Yang-Hui He, Romuald Janik, Matt Schwartz, Jesse Thaler, Dmitris Papailiopoulos, Jordan Ellenberg, Gary Shiu, and Yiqiao Zhong for fruitful discussions. We would like to thank Phil Wang (“lucidrains”) for use of his implementation of the RoPE and xPos relative position encodings. The work was supported in part by the U.S. Department of Energy (DOE) under Award No. DE-FOA-0002705, KA/OR55/22 (AIHEP). LD and TC are additionally supported by the U.S. Department of Energy Award No. DE-AC02-76SF00515. MW was supported by the research grant 00025445 from Villum Fonden.
@@ Appendix B Model Evaluation on the Full List of Linear Relations
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ Appendix C Coefficient-from-Key: Model Ablations
We perform a number of ablations in order to better understand the model performance for the coefficient-from-key experiment in Section 4, with an eye towards building a future model that can generalize across loops. Here we consider several different positional encoding schemes, explore encoder-only and decoder-only architectures, and vary both the number of layers in the model and the dimension $d$ of the encoded tokens. See Tables 5 and 7 for an overview of the results.
@@
C.1 Architecture Ablations: Relative Position Encoding, Encoder-only, Decoder-only, and Non-Transformer Architectures
@@
Finally, we perform the nonzero coefficient-from-key prediction task at 5 loops using non-Transformer architectures. Specifically, we use an encoder-decoder Long-Short-Term Memory recurrent neural network (LSTM) [45] and a Gated Recurrent Unit (GRU) [46]. The embedding dimension is $d=512$ and the number of layers in the encoder and decoder are both taken to be 2, as in the Transformer, while the hidden layer size is either chosen such that the number of trainable parameters is approximately the same as in the Transformer, or chosen to be $h=2048$. We again observe the familiar two-phase behavior as in Fig. 3. We report the results compared to a Transformer model in the table 6. For the models with a comparable number of parameters as the Transformer, we observe that performance is slightly worse. However, when $h=2048$, the LSTM performs at least as well as the Transformer on this particular task, and the GRU is not qualitatively worse.
Architecture
First Mag. Acc.
> 95% [Epoch]
Midpoint of
Step [Epoch]
First Total Acc.
> 95% [Epoch]
Best Acc.
Epoch 15
$N_{\text{params}}$
Transformer 3 6 9 99.10% 20M
LSTM ($h=786$) 7 12 N/A 87.42% 20M
GRU ($h=916$) 5 10 N/A 90.91% 20M
GRU ($h=2048$) 3 5 6 98.88% 85M
LSTM ($h=2048$) 3 5 6 99.76% 112M
Table 6: Model learning dynamics for a variety of LSTM architectures trained and evaluated on the same $L=5$ dataset. All display the double-plateau structure. We indicate the epoch at which the test-set magnitude accuracy first exceeds 95%, the midpoint of the plateau step, the epoch at which the overall test-set accuracy first exceeds 95%, and the best overall test-set accuracy after 15 epochs. The baseline Transformer model is indicated in bold. For this experiment, we use only one seed rather than the best of 4.
@@ C.2 Architecture Parameter Scan
Additionally, because we only split our data into training and test sets, rather than training, validation, and test sets, the accuracy metrics quoted may be unduly influenced by the specific choice of test set. In order to demonstrate that performance does not degrade appreciably for different choices of test set, we perform the 5-loop nonzero coefficient-from-key prediction task with three different random training/test set configurations, but the same architecture and hyperparameter configuration as in the baseline experiment in Section 4. We report these results in Table 8. The accuracies for the different splits are a little lower than for the baseline.
Architecture
First Mag. Acc.
> 95% [Epoch]
Midpoint of
Step [Epoch]
First Total Acc.
> 95% [Epoch]
Best Acc.
Epoch 15
Baseline 3 6 9 99.10%
Split 1 2 6 9 98.61%
Split 2 3 7 9 98.40%
Split 3 3 7 11 98.92%
Table 8: Model learning dynamics for four different choices of random training and test set split on the 5-loop coefficient-from-key task. We indicate the epoch at which the test-set magnitude accuracy first exceeds 95%, the midpoint of the plateau step, the epoch at which the overall test-set accuracy first exceeds 95%, and the best overall test-set accuracy after 15 epochs.